# K theoretic virtual fundamental cycle of global Kuranishi chart

### v1

@@ 2.2. Twisted sectors, inertia orbifold, and fake K-group
Here $h\in G$ acts on $(x,g)$ as $h\circ(x,g)=(hx,hgh^{-1})$, and $Y^{g}$ is the $g$-fixed locus in $Y$. We define $\mathcal{Y}^{(g)}$ (or simply $\mathcal{Y}^{g}$, if no confusion) as the component in $I\mathcal{Y}$ labeled by the conjugacy class $(g)$. We call $\mathcal{Y}^{(g)}$ sectors and call those with $g\neq 1$ twisted sectors.
@@
where $h\in G$ acts on $(x,\mathbf{g})$ by $h\circ(x,\mathbf{g})=(hx,h\mathbf{g}h^{-1})$. Define $\mathcal{Y}^{(\mathbf{g})}$ (or $\mathcal{Y}^{\mathbf{g}}$ if no confusion) as the sector in $I^{k}\mathcal{Y}$ corresponding to the conjugacy class $(\mathbf{g})$. We call $\mathcal{Y}^{(\mathbf{g})}$ a $k$-multi-sector.
By ‘fake’ theories of orbifolds, we mean theories in which we only remember the topological space structure on $\mathcal{Y}$ (and forget isotropy groups at points in $\mathcal{Y}$). As an example, define the fake orbifold K-group $K_{\mathbb{Z}}^{0,fake}(\mathcal{Y},\partial\mathcal{Y})$ as classes represented by $(\alpha=\alpha^{\ast}:\mathcal{E}\to\mathcal{E})\in K^{0}(\mathcal{Y},\partial\mathcal{Y})$, with isotropy groups at points in $\mathcal{Y}$ acting trivially on fibers in $\mathcal{E}$. Define
@@
Note that the super-trace of a $global$ isotropy $(g)\subseteq G$ defines a map $K^{0}(\mathcal{Y},\partial\mathcal{Y})\otimes\mathbb{C}\to K^{0,fake}(\mathcal{Y},\partial\mathcal{Y})$. Note also that by the eigen-bundle decomposition, there is an isomorphism
@@
by sending $\alpha$ to the element with $str_{(g)}i_{(g)}^{\ast}\alpha$ in the twisted sector labeled by $(g)$, where $i_{(g)}:\mathcal{Y}^{(g)}\to\mathcal{Y}$ is the inclusion.
@@ 2.3. K-homology
We consider Kasparov $\mathbb{C}(\mathcal{Y})$-modules $(H,\phi,T)$, where by definition $H$ is a $\mathbb{Z}_{2}$-graded complex Hilbert space, and $T$ a degree $1$ self-adjoint automorphism of $H$. The map $\phi:\mathbb{C}(\mathcal{Y})\to L(H)$ is a *-homomorphism from $\mathbb{C}(\mathcal{Y})$ to the space $L(H)$ of bounded linear maps on $H$, such that each $\phi(f)$ is of degree 00 and almost commutes with $T$ in the sense that $[T,\phi(f)]$ is compact in $L(H)$ [9]. The set of Kasparov $\mathbb{C}(\mathcal{Y})$-modules is an abelian semi-group under direct sum: $(H,\phi,T)\oplus(H^{\prime},\phi^{\prime},T^{\prime})=(H\oplus H^{\prime},\phi\oplus\phi^{\prime},T\oplus T^{\prime})$, with the zero element $H=0,\phi(f)=T=0$.
We say that $(H,\phi,T)$ is unitarily equivalent to $(H^{\prime},\phi^{\prime},T^{\prime})$, if there is a degree 00 unitary map $U:H\to H^{\prime}$ that takes $(\phi,T)$ to $(\phi^{\prime},T^{\prime})$. We say that $(H,\phi,T_{0})$ is homotopic to $(H,\phi,T_{1})$ if there is a continuous change of operators $T_{t}:t\in[0,1]$ that connects $T_{0},T_{1}$, such that $(H,\phi,T_{t})$ are Kasparov modules for all $t\in[0,1]$. We say that $(H,\phi,T)$ is degenerate if $T$ is invertible and commutes with $\phi(f)$ for all $f\in\mathbb{C}(\mathcal{Y})$.
We identify Kasparov modules that differ by a degenerated Kasparov $\mathbb{C}(\mathcal{Y}s)$-module and identify unitarily equivalent and homotopic Kasparov modules to obtain the fake $K$-homology (semi-)group $K^{fake}_{0}(\mathcal{Y})$. Indeed, $K^{fake}_{0}(\mathcal{Y})$ is a group [9]. The inverse of $(H,\phi,T)$ is $(H^{\prime},\phi^{\prime},-T)$, where $H^{\prime}$ is $H$ with the grading shifted by $1$, and $\phi^{\prime}=\phi\varepsilon$ with $\varepsilon$ acting as $(-1)^{a}$ on parts with grading $a$ (Proposition 1 in [9]).
@@
$\begin{array}[]{lll}f_{\ast}:&K^{fake}_{0}(\mathcal{Y})&\to K^{fake}_{0}(\mathcal{Y}^{\prime}).\\ &(H,\phi,T)&\to(H,\phi\circ f^{\ast},T)\end{array}$
@@ 2.4. Embedding relatively compact sets into compact orbifolds
In this subsection, we prove a technical proposition.
@@
Any compact subset $\mathcal{U}\subseteq\mathcal{Y}\setminus\partial\mathcal{Y}$ in a Riemannian orbifold $\mathcal{Y}$ has a neighborhood that open embeds into a closed orbifold $\mathcal{W}$.
@@
Assume without loss of generality that $\mathcal{U}$ is connected. By compactness of $\mathcal{U}$, there are finitely many geodesic balls $B(x_{i},r_{i})$ that cover $\mathcal{U}$, such that the exponential map at each $x_{i}$ is a diffeomorphism between $B(x_{i},2r_{i})$ and their pre-image. We start with an arbitrary compact co-dimension $1$ orbifold without boundary that meets some $B(x_{i},r_{i})$ and modify this orbifold on all $B(x_{i},r_{i})$’s to obtain a co-dimension $1$ orbifold that bounds a compact region $\mathcal{X}$ which contains all $B(x_{i},r_{i})$.
In each step of modification, we pick a geodesic ball $B(x_{i},r_{i})$ that meets the co-dimension $1$ orbifold and ‘push’ the sub-orbifold (black) out (red) of this ball within $B(x_{i},2r_{i})$, while fixing points near the boundary of $B(x_{i},2r_{i})$ (by pushing or pulling along the radius of corresponding geodesics emerging from $x_{i}$’s). Applying this construction until the resulting co-dimension $1$ orbifold $\partial\mathcal{X}$ does not meet any $B(x_{i},r_{i})$. In this case, $\partial\mathcal{X}$ bounds a compact region $\mathcal{X}$ that contains all $B(x_{i},r_{i})$, thus contains $\mathcal{U}$.
Moreover, since $\mathcal{U}$ is relatively compact, there are finitely many twisted strata. Under the assumption that the space $\mathcal{Y}$ is connected, each stratum is either the entire $\mathcal{Y}$ or a sub-orbifold of strictly lower dimension. We can make $\partial\mathcal{X}$ transverse to these strata by a perturbation. ∎
@@
There is a collar neighborhood $\varphi:\partial\mathcal{X}\times(0,1)\to\mathcal{X}$ of $\partial\mathcal{X}$ in $\mathcal{X}$. We take two copies $\mathcal{X}=\mathcal{X}^{\prime}$, with collar neighborhoods $\varphi,\varphi^{\prime}$, and define $\mathcal{W}$ as the ‘doubling’
@@ 2.5. The ‘choose-an-operator’ map
For a compact $G$-manifold $Y$ (with boundary), define $T^{\ast}_{G}Y$ as the set of co-vectors on $Y$ annihilating all vectors tangent to $G$-orbits. Add a $G$-invariant Hermitian structure on $Y$ and let $B^{\ast}_{G}Y$ be the unit ball in $T^{\ast}_{G}Y$. In this subsection, we construct a ‘choose-an-operator’ map $\mu=\mu_{G,Y}:K_{G}^{0,inv}(B_{G}^{\ast}Y,\partial B_{G}^{\ast}Y)\to K^{fake}_{0}(Y/G)$, where $K_{G}^{0,inv}$ stands for elements in $K_{G}^{0}$ represented by virtual bundles that correspond to the trivial representation of $G$. The orbifold ‘choose-an-operator’ map $\mu_{Y/G}$ is $\mu_{G,Y}$ sector-wise. The ‘choose-an-operator’ map when $G$ is the trivial group is constructed in, e.g., [5].
We first define certain ‘good’ representatives of elements in $K_{G}^{0,inv}(B_{G}^{\ast}Y,\partial B_{G}^{\ast}Y)$, which we call symbolic representatives, that serve as symbols of certain pseudo-differential operators representing K-homology classes.
If any two automorphisms of the same bundle agree outside a compact set, they represent the same element in the $K$ group. So to describe an element in $K_{G}^{0,inv}(B_{G}^{\ast}Y,\partial B_{G}^{\ast}Y)$, we only need to define the bundle involved and the automorphism near the boundary. Let $\pi_{Y}:B_{G}^{\ast}Y\to Y$, and let $p:\pi_{Y}^{\ast}E\to\pi_{Y}^{\ast}E$ be a bundle isomorphism, defined outside a compact subset which we refer to as $\operatorname{supp}p$, in $Y\setminus\partial Y\subseteq B_{G}^{\ast}Y$.
@@
A symbolic representative $p$ supported on $\operatorname{supp}p\subseteq B_{G}^{\ast}(Y)$ is an automorphism of $\pi_{Y}^{\ast}E\setminus E|_{\operatorname{supp}p}$, with
  • $E$ a $G$-bundle over $Y$ that corresponds to the trivial representation and is trivial outside $\operatorname{supp}p$
  • $p(y,\xi)=p(y,\nu\xi)$ for all $(y,\xi)\in B_{G}^{\ast}Y\setminus Y$ and $\nu\in\mathbb{R}_{+}$.
Symbolic representatives modulo certain relations (homotopy and isomorphism) comprise a group that is isomorphic to $K_{G}^{0,inv}(B_{G}^{\ast}Y,\partial B_{G}^{\ast}Y)$ (cf. [4] page 491-493, also [3] page 15). Moreover, we can take this symbolic representative to be self-adjoint.
By Corollary 2.1, we can equivariantly embed a $G$-invariant neighborhood of $\operatorname{supp}p$ into a closed manifold $W$ with almost free $G$-action. Let $E_{\pm}$ be the puls and minus part of the $\mathbb{Z}_{2}$ grading on $E$, and $p_{+}=p|_{E_{+}}$. There are sections of $E_{\pm}|_{Y\setminus\operatorname{supp}p}$ that serve as a basis on each fiber, under which the matrix of $p_{+}$ is the identity matrix. Thus we can extend $p,E$ trivially to $p_{W},E_{W}$ on $W$. Let $P_{W}:C^{\infty}(W,E_{W})\to C^{\infty}(W,E_{W})$ be the quantization of the (self-adjoint) symbolic representative $p_{W}$. Then $P_{W}$ is a self-adjoint pseudo-differential operator of order 00, and is of degree $1$ under the $\mathbb{Z}_{2}$-grading. If $s\in C^{\infty}(W,E_{W})$ is supported in $U\supseteq\operatorname{supp}p$, then $P_{W}s$ is also supported in $U$. If $s\in C^{\infty}(W,E_{W})$ is supported in $W\setminus\operatorname{supp}p$, then $P_{W}s=s$. Let $H=L^{2}(U,E|_{U})^{G}$ be the space of $G$-invariant $L^{2}$-sections of $E$ restricted to $U$, which is identified with a subset of $L^{2}(W,E_{W})$ by extending each section by 00 outside $U$. Since $P_{W}:C^{\infty}(W,E_{W})\to C^{\infty}(W,E_{W})$ is elliptic, it uniquely extends to $L^{2}$-sections in a continuous way. Let $T$ be the restriction to subspace $H$ of this continuous ($G$-equivariant) extension. Let ${\phi}$ be the $\mathbb{C}(Y/G)$-module structures on ${H}$, given by multiplying (the restriction to $U$ of) a continuous $G$-invariant function. We define
@@
Following [3], we show that $T$ is Fredholm by showing that the $G$-invariant parts $(\ker P_{W})^{G}=(\ker P_{W}^{\ast})^{G}$ are finite-dimensional. Pick a bi-invariant Riemannian metric on $G$, and let $v_{1},\cdots,v_{k}$ be first-order differential operators acting on sections of $E$ corresponding to some orthonormal basis of the Lie algebra of $G$. Consider the bundle map $\pi_{W}^{\ast}E\to\pi_{W}^{\ast}E$ given by
@@
The quantization of $\delta_{G}$ gives an order 00 pseudo-differential operator whose kernel contains the space of $G$-invariant functions as a closed subspace. Now, the operator $A=(P_{W},\Delta_{G})$ has an injective symbol, so the symbol of $A^{\ast}A$ is invertible. Thus the operator $A^{\ast}A=P_{W}^{\ast}P_{W}+\Delta^{\ast}_{G}\Delta_{G}$ is elliptic with finite-dimensional kernel. Now $(\ker P_{W})^{G}=(\ker P^{\ast}_{W})^{G}$ is contained in the kernel of $A^{\ast}A$, so it is also finite-dimensional. Thus $T$ is Fredholm.
Moreover, for an arbitrary $f$, the order 00 symbol of the commutator $[P_{W}|_{U},\phi(f)]$ is 00, so $[T,\phi(f)]$ is compact. Thus $\mu_{G,Y}(\pi_{W}^{\ast}E_{W},p_{W})\in K^{fake}_{0}(Y)$. ∎
@@
The operator $\mu_{G,Y}$ is well-defined as a map $K_{G}^{0,inv}(B_{G}^{\ast}Y,\partial B_{G}^{\ast}Y)\to K_{0}^{fake}(Y)$.
@@
We temporarily denote the ‘choose-an-operator’ map with choice $U,W,p$ by $\mu_{G,Y}^{U,W}(p)$. We need to prove:
@@
    $\mu_{G,Y}^{U,W}([p])$ is independent of $U,W$.
For fixed $U,W$, let $P_{W},P_{W}^{\prime}$ be two pseudo-differential operators with symbolic representatives $[p_{W}]=[p_{W}^{\prime}]$. If $p_{W},p_{W}^{\prime}$ are homotopic (in particular, if $p_{W}=p_{W}^{\prime}$ but $P_{W}\neq P_{W}^{\prime}$), i.e., their symbols are connected by a homotopy of degree $1$ automorphisms of the $same$ bundle $\pi^{\ast}_{W}E$, then the Hilbert spaces involved are the same, and the operators are connected by a homotopy of operators given by the quantization of the symbols, thus connected by a homotopy of representatives of $K$-homology. If the $p_{W}$ is degenerated $K_{G}^{0,inv}(B_{G}^{\ast}W,\partial B_{G}^{\ast}W)$, then it is an isomorphism between positive and negative parts in the $\mathbb{Z}_{2}$-grading of the ambient bundle with empty support. Computing under the identification given by this isomorphism, the operator $T$ is identity, and thus $\mu_{G,Y}(p_{W})$ is degenerated.
For fixed $U$, the Hilbert spaces are naturally identified, and different $W$’s give the same operator $T$. For $\operatorname{supp}p\subseteq U^{\prime}\subseteq U$, let $(E,p)\in K_{G}^{0,inv}(B_{G}^{\ast}Y,\partial B_{G}^{\ast}Y)$ be a symbolic representative. The corresponding $K$-homology element defined by $U^{\prime}$ and $U$ differs by the image of the degenerated element $(E,p)|_{U\setminus U^{\prime}}$ under the choose-an-operator map, thus they define the same element in $K_{0}^{fake}(Y)$. Two general compact sets are contained in their union, so they also define the same element in $K_{0}^{fake}(Y)$. ∎
@@
The cotangent bundle of $\mathcal{Y}=Y/G$ is $T^{\ast}\mathcal{Y}=(T^{\ast}_{G}Y)/G$. Denote its unit ball bundle by $B^{\ast}\mathcal{Y}$ (under a Hermitian metric). Observe that the sector $(B^{\ast}\mathcal{Y})^{(g)}=B^{\ast}(\mathcal{Y})^{(g)}$, and that the supertrace $str_{g}$ maps $K^{0,fake}(B^{\ast}\mathcal{Y}^{(g)},\partial B^{\ast}\mathcal{Y}^{(g)})$ to $K_{G}^{0,inv}(B_{G}^{\ast}Y^{(g)},\partial B_{G}^{\ast}Y^{(g)})\otimes\mathbb{C}$. We define the ‘genuine’ choose-an-operator map $\mu$ sector-wise, as the invariant choose-an-operator map composed with supertraces. The orbifold choose-an-operator map is independent of the presentation of the orbifold.

### v2

@@ 2.2. Twisted sectors, inertia orbispace, and fake K-group
Here $h\in G$ acts on $(x,g)$ as $h\circ(x,g)=(hx,hgh^{-1})$, and $Y^{g}$ is the $g$-fixed locus in $Y$. We define $\mathcal{Y}^{(g)}$ (or simply $\mathcal{Y}^{g}$, when no confusion arises) as the component in $I\mathcal{Y}$ labeled by the conjugacy class $(g)$. We call the $\mathcal{Y}^{(g)}$ the sectors, and those with $g\neq 1$ the twisted sectors.
@@
where $h\in G$ acts on $(x,\mathbf{g})$ by $h\circ(x,\mathbf{g})=(hx,h\mathbf{g}h^{-1})$. Define $\mathcal{Y}^{(\mathbf{g})}$ (or $\mathcal{Y}^{\mathbf{g}}$ when no confusion arises) as the sector in $I^{k}\mathcal{Y}$ corresponding to the conjugacy class $(\mathbf{g})$. We call $\mathcal{Y}^{(\mathbf{g})}$ a $k$-multi-sector.
By ‘fake’ theories of orbispaces, we mean theories in which we only remember the topological space structure on $\mathcal{Y}$ (and forget the isotropy groups at points in $\mathcal{Y}$). As an example, define the fake orbispace K-group $K_{\mathbb{Z}}^{0,fake}(\mathcal{Y},\partial\mathcal{Y})$ as classes represented by $(\alpha=\alpha^{\ast}:\mathcal{E}\to\mathcal{E})\in K^{0}(\mathcal{Y},\partial\mathcal{Y})$, with the isotropy groups at points in $\mathcal{Y}$ acting trivially on the fibers of $\mathcal{E}$. Define
@@
Note that the super-trace of a $global$ isotropy $(g)\subseteq G$ defines a map
$K^{0}(\mathcal{Y},\partial\mathcal{Y})\otimes\mathbb{C}\to K^{0,fake}(\mathcal{Y},\partial\mathcal{Y}).$
Note also that by the eigen-bundle decomposition, there is an isomorphism
@@
given by sending $\alpha$ to the element with $str_{(g)}\,i_{(g)}^{\ast}\alpha$ in the twisted sector labeled by $(g)$, where $i_{(g)}:\mathcal{Y}^{(g)}\to\mathcal{Y}$ is the inclusion.
@@ 2.3. K-homology
We consider Kasparov $\mathbb{C}(\mathcal{Y})$-modules $(H,\phi,T)$, where by definition $H$ is a $\mathbb{Z}_{2}$-graded complex Hilbert space and $T$ is a degree $1$ self-adjoint automorphism of $H$. The map $\phi:\mathbb{C}(\mathcal{Y})\to L(H)$ is a *-homomorphism from $\mathbb{C}(\mathcal{Y})$ to the space $L(H)$ of bounded linear maps on $H$, such that each $\phi(f)$ is of degree 00 and almost commutes with $T$ in the sense that $[T,\phi(f)]$ is compact in $L(H)$ [9]. The set of Kasparov $\mathbb{C}(\mathcal{Y})$-modules is an abelian semi-group under direct sum: $(H,\phi,T)\oplus(H^{\prime},\phi^{\prime},T^{\prime})=(H\oplus H^{\prime},\phi\oplus\phi^{\prime},T\oplus T^{\prime})$, with the zero element defined by $H=0$, $\phi(f)=0$, and $T=0$.
We say that $(H,\phi,T)$ is unitarily equivalent to $(H^{\prime},\phi^{\prime},T^{\prime})$ if there is a degree 00 unitary map $U:H\to H^{\prime}$ that takes $(\phi,T)$ to $(\phi^{\prime},T^{\prime})$. We say that $(H,\phi,T_{0})$ is homotopic to $(H,\phi,T_{1})$ if there exists a continuous family of operators $\{T_{t}\}_{t\in[0,1]}$ connecting $T_{0}$ and $T_{1}$ such that $(H,\phi,T_{t})$ is a Kasparov module for all $t\in[0,1]$. We say that $(H,\phi,T)$ is degenerate if $T$ is invertible and commutes with $\phi(f)$ for all $f\in\mathbb{C}(\mathcal{Y})$.
We identify Kasparov modules that differ by a degenerate Kasparov $\mathbb{C}(\mathcal{Y})$-module and also identify unitarily equivalent and homotopic Kasparov modules to obtain the fake $K$-homology (semi-)group $K^{fake}_{0}(\mathcal{Y})$. Indeed, $K^{fake}_{0}(\mathcal{Y})$ is a group [9]. The inverse of $(H,\phi,T)$ is $(H^{\prime},\phi^{\prime},-T)$, where $H^{\prime}$ is $H$ with the grading shifted by $1$, and $\phi^{\prime}=\phi\varepsilon$ with $\varepsilon$ acting as $(-1)^{a}$ on components of grading $a$ (Proposition 1 in [9]).
@@
$\begin{array}[]{lll}f_{\ast}:&K^{fake}_{0}(\mathcal{Y})&\to K^{fake}_{0}(\mathcal{Y}^{\prime}),\\ &(H,\phi,T)&\to(H,\phi\circ f^{\ast},T).\end{array}$
@@ 2.4. Embedding relatively compact sets into compact orbifolds
Starting from this subsection, we assume $\mathcal{Y}$ is an orbifold. In this subsection, we prove a technical proposition.
@@
Any compact subset $\mathcal{U}\subseteq\mathcal{Y}\setminus\partial\mathcal{Y}$ in a Riemannian orbifold $\mathcal{Y}$ has a neighborhood that embeds openly into a closed orbifold $\mathcal{W}$.
@@
Assume without loss of generality that $\mathcal{U}$ is connected. By the compactness of $\mathcal{U}$, there are finitely many geodesic balls $B(x_{i},r_{i})$ that cover $\mathcal{U}$, such that the exponential map at each $x_{i}$ is a diffeomorphism between $B(x_{i},2r_{i})$ and its pre-image. We start with an arbitrary compact codimension-$1$ orbifold without boundary that meets some $B(x_{i},r_{i})$ and modify this orbifold on all the $B(x_{i},r_{i})$’s to obtain a codimension-$1$ orbifold that bounds a compact region $\mathcal{X}$ which contains all the $B(x_{i},r_{i})$.
In each step of the modification, we pick a geodesic ball $B(x_{i},r_{i})$ that meets the codimension-$1$ orbifold and “push” the sub-orbifold (black) out (red) of this ball within $B(x_{i},2r_{i})$, while fixing points near the boundary of $B(x_{i},2r_{i})$ (by pushing or pulling along the radii of the corresponding geodesics emerging from the $x_{i}$’s). We apply this construction until the resulting codimension-$1$ orbifold $\partial\mathcal{X}$ does not meet any $B(x_{i},r_{i})$. In this case, $\partial\mathcal{X}$ bounds a compact region $\mathcal{X}$ that contains all the $B(x_{i},r_{i})$, and hence contains $\mathcal{U}$.
Moreover, since $\mathcal{U}$ is relatively compact, there are only finitely many twisted strata. Under the assumption that the space $\mathcal{Y}$ is connected, each stratum is either the entire $\mathcal{Y}$ or a sub-orbifold of strictly lower dimension. We can make $\partial\mathcal{X}$ transverse to these strata by a perturbation. ∎
@@
There is a collar neighborhood $\varphi:\partial\mathcal{X}\times(0,1)\to\mathcal{X}$ of $\partial\mathcal{X}$ in $\mathcal{X}$. We take two copies $\mathcal{X}=\mathcal{X}^{\prime}$, with collar neighborhoods $\varphi$ and $\varphi^{\prime}$, and define $\mathcal{W}$ as the “doubling”
@@ 2.5. The ‘choose-an-operator’ map
For a compact $G$-manifold $Y$ (with boundary), define $T^{\ast}_{G}Y$ as the set of covectors on $Y$ that annihilate all vectors tangent to $G$-orbits. Equip $Y$ with a $G$-invariant Hermitian structure and let $B^{\ast}_{G}Y$ be the unit ball in $T^{\ast}_{G}Y$. In this subsection, we construct a ‘choose-an-operator’ map $\mu=\mu_{G,Y}:K_{G}^{0,inv}(B_{G}^{\ast}Y,\partial B_{G}^{\ast}Y)\to K^{fake}_{0}(Y/G)$, where $K_{G}^{0,inv}$ stands for those elements in $K_{G}^{0}$ represented by virtual bundles corresponding to the trivial representation of $G$. The orbifold ‘choose-an-operator’ map $\mu_{Y/G}$ is defined sector-wise via $\mu_{G,Y}$. (The ‘choose-an-operator’ map when $G$ is the trivial group is constructed in, e.g., [5].)
We first define certain “good” representatives of elements in $K_{G}^{0,inv}(B_{G}^{\ast}Y,\partial B_{G}^{\ast}Y)$, which we call symbolic representatives, that serve as the symbols of certain pseudo-differential operators representing K-homology classes.
If any two automorphisms of the same bundle agree outside a compact set, they represent the same element in the $K$-group. Thus, to describe an element in $K_{G}^{0,inv}(B_{G}^{\ast}Y,\partial B_{G}^{\ast}Y)$, we only need to define the bundle involved and the automorphism near the boundary. Let $\pi_{Y}:B_{G}^{\ast}Y\to Y$, and let $p:\pi_{Y}^{\ast}E\to\pi_{Y}^{\ast}E$ be a bundle isomorphism defined outside a compact subset—which we refer to as $\operatorname{supp}p$—in $Y\setminus\partial Y\subseteq B_{G}^{\ast}Y$.
@@
A symbolic representative $p$ supported on $\operatorname{supp}p\subseteq B_{G}^{\ast}(Y)$ is an automorphism of $\pi_{Y}^{\ast}E$ outside $E|_{\operatorname{supp}p}$, with
  • $E$ a $G$-bundle over $Y$ that corresponds to the trivial representation and is trivial outside $\operatorname{supp}p$,
  • $p(y,\xi)=p(y,\nu\xi)$ for all $(y,\xi)\in B_{G}^{\ast}Y\setminus Y$ and for all $\nu\in\mathbb{R}_{+}$.
Symbolic representatives, modulo the relations of homotopy and isomorphism, form a group that is isomorphic to $K_{G}^{0,inv}(B_{G}^{\ast}Y,\partial B_{G}^{\ast}Y)$ (cf. [4] page 491-493, also [3] page 15). Moreover, we can take this symbolic representative to be self-adjoint.
By Corollary 2.1, we can equivariantly embed a $G$-invariant neighborhood of $\operatorname{supp}p$ into a closed manifold $W$ with an almost free $G$-action. Let $E_{\pm}$ be the positive and negative parts of the $\mathbb{Z}_{2}$-grading on $E$, and set $p_{+}=p|_{E_{+}}$. There exist sections of $E_{\pm}|_{Y\setminus\operatorname{supp}p}$ that form a basis on each fiber, under which the matrix of $p_{+}$ is the identity. Thus, we can extend $p$ and $E$ trivially to $p_{W}$ and $E_{W}$ on $W$. Let $P_{W}:C^{\infty}(W,E_{W})\to C^{\infty}(W,E_{W})$ be the quantization of the (self-adjoint) symbolic representative $p_{W}$. Then $P_{W}$ is a self-adjoint pseudo-differential operator of order 00 and is of degree $1$ under the $\mathbb{Z}_{2}$-grading. If $s\in C^{\infty}(W,E_{W})$ is supported in some $U\supseteq\operatorname{supp}p$, then $P_{W}s$ is also supported in $U$. If $s\in C^{\infty}(W,E_{W})$ is supported in $W\setminus\operatorname{supp}p$, then $P_{W}s=\psi s$ for some invertible map $\psi$ on $W\setminus\operatorname{supp}p$. Let $H=L^{2}(U,E|_{U})^{G}$ be the space of $G$-invariant $L^{2}$-sections of $E$ restricted to $U$, which is identified with a subspace of $L^{2}(W,E_{W})$ by extending each section by 00 outside $U$. Since $P_{W}:C^{\infty}(W,E_{W})\to C^{\infty}(W,E_{W})$ is elliptic, it uniquely extends in a continuous way to $L^{2}$-sections. Let $T$ be the restriction of this continuous ($G$-equivariant) extension to the subspace $H$. Let $\phi$ denote the $\mathbb{C}(Y/G)$-module structure on $H$, given by multiplication (restricted to $U$) by a continuous $G$-invariant function. We define
@@
Following [3], we show that $T$ is Fredholm by proving that the $G$-invariant parts $(\ker P_{W})^{G}=(\ker P_{W}^{\ast})^{G}$ are finite-dimensional. Pick a bi-invariant Riemannian metric on $G$, and let $v_{1},\cdots,v_{k}$ be first-order differential operators acting on sections of $E$ corresponding to an orthonormal basis of the Lie algebra of $G$. Consider the bundle map $\pi_{W}^{\ast}E\to\pi_{W}^{\ast}E$ given by
@@
The quantization of $\delta_{G}$ yields an order 00 pseudo-differential operator whose kernel contains the space of $G$-invariant functions as a closed subspace. Now, the operator $A=(P_{W},\Delta_{G})$ has an injective symbol, so the symbol of $A^{\ast}A$ is invertible. Thus, the operator $A^{\ast}A=P_{W}^{\ast}P_{W}+\Delta^{\ast}_{G}\Delta_{G}$ is elliptic with finite-dimensional kernel. Since $(\ker P_{W})^{G}=(\ker P_{W}^{\ast})^{G}$ is contained in the kernel of $A^{\ast}A$, it is also finite-dimensional. Hence, $T$ is Fredholm.
Moreover, for an arbitrary $f$, the order 00 symbol of the commutator $[P_{W}|_{U},\phi(f)]$ is zero, so $[T,\phi(f)]$ is compact. Thus, $\mu_{G,Y}(\pi_{Y}^{\ast}E,p)$ represents an element in $K^{fake}_{0}(Y)$. ∎
@@
The operator $\mu_{G,Y}$ is well-defined as a map
$K_{G}^{0,inv}(B_{G}^{\ast}Y,\partial B_{G}^{\ast}Y)\to K_{0}^{fake}(Y).$
@@
We temporarily denote the ‘choose-an-operator’ map with the choices $U,W,p$ by $\mu_{G,Y}^{U,W}(p)$. We need to prove:
@@
    $\mu_{G,Y}^{U,W}([p])$ is independent of the choices of $U$ and $W$.
For fixed $U$ and $W$, let $P_{W}$ and $P_{W}^{\prime}$ be two pseudo-differential operators with symbolic representatives satisfying $[p_{W}]=[p_{W}^{\prime}]$. If $p_{W}$ and $p_{W}^{\prime}$ are homotopic (in particular, if $p_{W}=p_{W}^{\prime}$ but $P_{W}\neq P_{W}^{\prime}$), i.e., if their symbols are connected by a homotopy of degree $1$ automorphisms of the same bundle $\pi^{\ast}_{W}E$, then the Hilbert spaces involved are the same and the operators are connected by a homotopy (given by the quantization of the symbols), hence representing the same class in $K$-homology. If $p_{W}$ is degenerate in $K_{G}^{0,inv}(B_{G}^{\ast}W,\partial B_{G}^{\ast}W)$, then it is an isomorphism between the positive and negative parts in the $\mathbb{Z}_{2}$-grading of the ambient bundle with empty support. Under the identification given by this isomorphism, the operator $T$ is the identity, and thus $\mu_{G,Y}(p_{W})$ is degenerate.
For fixed $U$, the Hilbert spaces are naturally identified, and different choices of $W$ yield the same operator $T$. For $\operatorname{supp}p\subseteq U^{\prime}\subseteq U$, let $(E,p)\in K_{G}^{0,inv}(B_{G}^{\ast}Y,\partial B_{G}^{\ast}Y)$ be a symbolic representative. The corresponding $K$-homology element defined using $U^{\prime}$ versus $U$ differs by the image of the degenerate element $(E,p)|_{U\setminus U^{\prime}}$ under the choose-an-operator map. Thus, they define the same element in $K_{0}^{fake}(Y)$. Since any two general compact sets are contained in their union, the defined element in $K_{0}^{fake}(Y)$ is independent of the choices of $U$ and $W$. ∎
@@
The cotangent bundle of $\mathcal{Y}=Y/G$ is $T^{\ast}\mathcal{Y}=(T^{\ast}_{G}Y)/G$. Denote its unit ball bundle by $B^{\ast}\mathcal{Y}$ (with respect to a Hermitian metric). Observe that the sector $(B^{\ast}\mathcal{Y})^{(g)}=B^{\ast}(\mathcal{Y})^{(g)}$, and that the supertrace $str_{g}$ maps $K^{0,fake}(B^{\ast}\mathcal{Y}^{(g)},\partial B^{\ast}\mathcal{Y}^{(g)})$ to $K_{G}^{0,inv}(B_{G}^{\ast}Y^{(g)},\partial B_{G}^{\ast}Y^{(g)})\otimes\mathbb{C}$. We define the ‘genuine’ choose-an-operator map $\mu$ sector-wise as the composition of the invariant choose-an-operator map with the supertraces. The orbifold choose-an-operator map is independent of the presentation of the orbifold.